16. Calculate volume ratio vi for an adiabatic change (Pi,V) → (P,-2 × P1%) of one mole of HsO gas. The gas can be treated as an ideal gas.
Answer #1

Given :

At state 1

Pressure = P1

Volume = V1

At state 2

Pressure = P2 = 2P1

Volume = V 2 , H2O is treated as ideal gas

For an adiabatic process

PV^(\gamma ) = constant.........(1)

where \gamma = Heat capacity ratio ( Cp/ Cv)

From eqn 1 , we can write as

P1 *V1^(\gamma) = P2 * V2^( \gamma )

[V1/V2 ]=[ P2/ P1]^(1/\gamma)

V1/ V2 = [ 2 ] ^(1/1.33) { given : P2/ P1= 2 }

( Note: For tri- atomic gas , H2O, \gamma = 1.33 )

V1/V2 = 1.683

Hence volume ratio, V1/ V2 = 1.683